# Cedegao E. Zhang1,*

### v1

@@ Abstract
We can evaluate features of problems and their potential solutions well before we can effectively solve them. When considering a game we have never played, for instance, we might infer whether it is likely to be challenging, fair, or fun simply from hearing the game rules, prior to deciding whether to invest time in learning the game or trying to play it well. Many studies of game play have focused on optimality and expertise, characterizing how people and computational models play based on moderate to extensive search and after playing a game dozens (if not thousands or millions) of times. Here, we study how people reason about a range of simple but novel connect-n style board games. We ask people to judge how fair and how fun the games are from very little experience: just thinking about the game for a minute or so, before they have ever actually played with anyone else, and we propose a resource-limited model that captures their judgments using only a small number of partial game simulations and almost no lookahead search.
@@ Introduction
Figure 1: (A) Design of 121 novel grid games varying the game environment, dynamics, and win conditions. (B) Our intuitive game theory model simulates bounded game play under a fast but general agent model to draw probabilistic inferences about novel games.
The formal and empirical study of games often focuses on optimality and expertise. Game theory in mathematical economics, as studied by \@BBOPcite\@BAP\@BBNNash (1950)\@BBCP for example, largely seeks to characterize equilibrium states reached under rational play. Computational game playing models, from classic AI expert systems like Deep Blue to more recent neurally guided models like AlphaGo and AlphaZero \@BBOPcitep\@BAP\@BBN(Silver et al., 2016, 2017, 2018)\@BBCP, aim to match or exceed the abilities of human experts, generally by evaluating millions of simulated games. A rich body of related cognitive science research documents how human experts play after extensive experience on a given game \@BBOPcitep\@BAP\@BBN(Newell & Simon, 1972; Gobet et al., 2004)\@BBCP.
Experts seem to learn highly specialized and game-specific representations, like memorized chess states and openings, that enable them to effectively search or approximate the results of searching much further ahead in a game than novice players \@BBOPcitep\@BAP\@BBN(Chase & Simon, 1973; van Opheusden et al., 2023)\@BBCP.
@@
The present study focuses on this higher-level capacity to flexibly evaluate novel inference and decision making problems, using a set of novel strategy games that vary the environment, dynamics, and win conditions of more familiar grid games as a test bed for general problem evaluation. We propose a computational model (Fig. 1b) that implements novice but goal-directed agents as fast but search-limited planners. Our model estimates the value of intermediate game states based on simple but general concepts about game play, such as making progress towards a sparse overall goal (like making $M$ in a row) and preventing opponent progress when in a competitive setting. We then model how people evaluate games overall by nesting this agent model within a sample-based inference procedure that estimates game outcomes from a limited number of partial game simulations. While we focus here on games, the underlying building blocks of this “intuitive game theory” model are designed to capture more basic multi-agent planning and probabilistic inference components applicable to a much broader range of problems—where people are called on to make good guesses and good bets in novel situations, and to evaluate their prospects of success along with the emotional reward of engaging \@BBOPcitep\@BAP\@BBN(Allen et al., 2024)\@BBCP.
@@
We find that our intuitive game theory model, despite using only a single step of search in its agent model and a small number of partial game simulations to estimate game play, well predicts participants’ judgements of game outcomes on novel games. We compare our model against several alternative models, which either (i) consider even more naive game agent models, or conversely (ii) expend much more computational effort on search and game simulation to more optimally predict outcomes. We also compare against two language-model-based baselines with no explicit agent model that reason solely from background knowledge about other more familiar games.
Using our model, we also show that several game-related factors correlate with human fun ratings on these games: a fairness factor based on estimated entropy over game outcomes; a challengingness factor based on the estimated advantage our model has in playing against a random agent; and a fun factor based on language model predictions. Combined into a joint regression model, these factors together can predict most of the variance in human fun judgements.
Category Explanation Count
@@
To choose any single action, an optimal agent model should estimate and then seek to maximize downstream utility with respect to all available actions from the current game state. Many general game play models estimate utility for novel games using expensive lookahead search, ideally to terminal reward-generating game states \@BBOPcitep\@BAP\@BBN(Genesereth & Thielscher, 2014; Yannakakis & Togelius, 2018)\@BBCP. For instance, the class of games we consider here defines very sparse utility functions—winning or losing is only determined based on a final state in which a player or their opponent has placed $M$ pieces in a row. In cases where $M$ is large or players are playing on large boards (e.g., the goal is to make 8 in a row on a 20 by 20 board), this means searching over potentially long trajectories of game play to evaluate any single action.
To then draw probabilistic judgements about game outcomes (such as how likely a given non-deterministic player is to win on average), generic sample-based inference based on this agent model would require running multiple full game simulations, wherein each agent action would require downstream search to these terminal states.
In contrast, our intuitive game theory model aims to produce fast, probabilistic inferences over a broad range of novel games. This model simulates agent game play based on basic but general assumptions about goal-oriented players that can be estimated directly from intermediate states, then draws graded judgements based on partial game simulations over a range of game depths rather than always simulating to terminal states.
@@
The core of our model is a simulated agent that can play from arbitrary game specifications. This agent model has a general search-limited way of placing utility over intermediate states.
@@
Specifically, given an intermediate game state, our player model assigns an overall utility score to each possible next move (each open position on the board) based on the following functions:
@@
    The maximum number of connected contiguous pieces that the position entails on any of the allowed winning directions (horizontal, vertical, or diagonal in general—modulated with respect to any player-specific restrictions on win directions), $n_{1}\in\mathbb{Z}^{+}$. If $n_{1}$ equals the winning $m$ for $m$ in a row, we add an additional 1 to $n_{1}$. This function captures the general intuition that players are goal-directed and try to make progress towards winning—that is, that players derive utility from intermediate states towards the most rewarding terminal state. While there are many possible subgoals that a player could consider which might make partial progress towards terminal states (e.g., placing making 3 in a row by placing two unconnected pieces and then placing one more piece in the center), we choose a simple, easy to calculate function that also follows from players following an intuitive standard policy (making $m$ in a row by previously building up to $m-1$ in a row, $m-2$ in a row, and so on via contiguous pieces). We also only consider the maximum number of pieces to reflect a relatively naive, shallow player who is considering a single, simple policy towards the nearest victory. This function could easily be augmented to support a number of related functions that reflect partial progress towards multiple terminal rewarding states.
@@
    Relatedly, the maximum number of connected contiguous pieces that the position blocks the opponent from having, $n_{2}\in\mathbb{Q}^{+}$, also considered with respect to the win-directions applicable to the opponent. This function captures the intuition that competitive players in antagonistic two-player games should also attempt to block opponent progress towards opponent goals. Note that this function is simple, symmetric to our definition of player progress (because players in our games have related and opposing goals). It also effectively evaluates just a single step of lookahead search relative to a opponent who is also relatively naive (that is, an opponent who at the next step would simply attempt to make progress towards their nearest victory along a simple policy). We subtract $0.5$ from $n_{2}$ to reflect people’s tendency to weigh offense over defense \@BBOPcitep\@BAP\@BBN(Crowley & Siegler, 1993)\@BBCP, so blocking the opponent’s $\hat{n}$ in a row is not as good as making an $\hat{n}$ in a row for oneself (but is better than making an $\hat{n}-1$ in a row). However, if $n_{2}$ equals the winning $m$, we do not subtract $0.5$. One could consider other related or more complex formulations that also capture the notion of blocking opponent progress.
These components together reflect the intuitions that people often but not always place pieces around the center at the beginning, people naturally try to make progress towards winning, and people know they should secure winning moves and block the opponent from having those. The utility of a position $p$ is then given by an exponential function:
$V(p)=2^{((1-d)+n_{1}+n_{2})}.$
We apply a softmax (temperature = $1$) to the set of position utilities to obtain a probability distribution over possible moves, and the agent samples a move from this distribution.
Given this model, how do we draw graded inferences about how fair a game is? For each given game, the model runs $k$ simulations of two players playing the game. Here we set $k=20$, building upon the idea that people use a relatively small number of simulations to form beliefs and make decisions \@BBOPcitep\@BAP\@BBN(Vul et al., 2014; Icard, 2016)\@BBCP. The model does not necessarily simulate until a terminal state of the game (i.e., a win, loss, or draw happens). Instead, the model randomly samples a number of moves it would play out from $1$ to the size of the game board, and evaluates the game outcome based on the board state only up to that maximum number of moves (if a terminal win state hasn’t been reached, we evaluate the board as a draw). This reflects the notion that one could stop a simulation after making a few moves, stop in the middle, or play until the end, and such behaviors are observed in the human scratchpad usage data. Given the numbers of win, loss, and draw outcomes in these $20$ move-capped simulations, we can compute the expected payoff for the game (with scores $1$ for win, $-1$ for loss, and 00 for draw).
@@
Figure 3: Expected payoff across models against humans-predicted payoff. Each point represents the payoff for one of the $n=121$ game stimuli. Error bars in the y-axis are over individual human judgments per game and depict standard error.
@@
Participants We recruit $484$ participants from Prolific \@BBOPcitep\@BAP\@BBN(Palan & Schitter, 2018)\@BBCP. Each participant was randomly presented with $10$ games sampled from our stimuli, as well as regular Tic-Tac-Toe (won by making 3 in a row on a $3\times 3$ board) to normalize game judgements. We collect approximately $20$ judgements per game stimuli for each game reasoning query. Participants were paid at a base rate of \$12.5/hr with an optional bonus up to \$15/hr; the full experiment approximately took $25$ minutes.
@@
Game outcome and game fun judgments Subjects were randomly assigned to one of two game reasoning conditions. In the game outcome evaluation condition, subjects produced judgements on a continuous 0-100 probability scale to predict the likelihood of a first player win (if the game does not end in a draw, assuming both players play reasonably, how likely is it that the first player is going to win (not draw)?) and a draw (assuming both players play reasonably, how likely is the game to end in a draw?). In the game fun condition, subjects instead assessed the likelihood that the game is fun (how fun is this game?) on a confidence scale spanning 0 (the least fun of this class of game) to 100 (the most fun of this class of game).
@@
Participants produced judgements about each game based on a linguistic game specification. We additionally provided participants with an interactive scratchpad board that they were told they could, but were not required, to use to inform their judgements; the scratchpad permitted automatically placing pieces of different colors (to simulate players) and could be cleared entirely to begin a new game. Participants were required to consider each game for at least $60$ seconds before producing game judgements; on average, participants took 87.2 ($\pm$ 2.5 STE) seconds per game in the game outcome condition and 83.4 ($\pm$ 2.41 STE) seconds per game in the fun rating condition.
@@
Figure 4: Absolute difference between human- and model-predicted payoff, broken down by game category. Error bars depict 95% CI over games within each game category (number of games in each game category is depicted in 1). The MCS-based model has an average distance over 0.7 on nearly all games and is omitted from these plots.
@@
Novel game creation After answering all game reasoning queries, we additionally asked participants to create a new grid-based game variant that they would find fun. Participants wrote a linguistic game specification, describing the board size and win conditions. As in the game judgement queries, participants were again provided an optional scratchpad and required to spend $60$ seconds before submitting a response. The scratchpad enabled participants to try out the game they intend to create. After specifying a game, participants were asked to answer the same game reasoning query (either game outcomes or game fun) about their own game.
Alternative computational models
@@
We implement a range of alternative computational models designed to span both weaker estimates of game play (using more naive agent models, or without any explicit agent models) and more optimal estimates of game play (using additional search and game simulations to terminal states).
@@
Partial game simulations with random agent This baseline estimates outcomes using partial game simulations, but substitutes our sub-goaling agent model with a more naive (“random”) agent model that selects actions from a uniform distribution over valid moves.
@@
Full game simulations with random agent This baseline infers outcomes using complete game simulations to terminal states, but using the random agent model.
Full game simulations with depth-5 lookahead search This baseline infers outcomes using complete game simulations to terminal states, and chooses moves based on future state utilities (under our sub-goaling utility metric) using a depth-5 lookahead search (i.e., we model alternating agent play using our model up to 5 actions ahead, then choose actions based on utility estimates under our subgoaling function at the future state). The search-based agent model is inspired by the finding in \@BBOPcite\@BAP\@BBNvan Opheusden et al. (2023)\@BBCP, which suggests that experts playing a specific Connect-4 grid game are well-modeled by depth-5 search.
Full game simulations with Monte Carlo Search (MCS) This baseline infers outcomes using complete game simulations to terminal states. The agent employs a standard MCS utility estimator \@BBOPcitep\@BAP\@BBN(Genesereth & Thielscher, 2014)\@BBCP, which runs full game simulations under a random agent model to estimate action utilities and then selects moves from the softmax distribution over these estimates. MCS is often used to approximate oracle, near-optimal game utility estimates with many simulations; we use $k=20$ simulations per move.
Large Language Model (LLM) game evaluations We also implement two LLM baselines to evaluate game reasoning with no explicit agent model, but rather, models which can draw on background knowledge about classic grid games and variants (including extensive information, for instance, about the optimal solutions to games like Tic-Tac-Toe and Connect-4). We use GPT-4 \@BBOPcitep\@BAP\@BBN(OpenAI, 2023)\@BBCP as our base LLM and implement two LLM-based game evaluation baselines: a zero-shot baseline (which is directly queried with the linguistic game specification and game reasoning query), and a zero-shot chain-of-thought (CoT) baseline \@BBOPcitep\@BAP\@BBN(Kojima et al., 2022; Wei et al., 2022)\@BBCP, which is first prompted to produce additional reasoning in text before answering a query (a standard method used to account for the computational limits in directly producing answers to questions given the LLM architecture). In both baselines, we prompt the model with a lightly-modified variant of the full experiment instructions shown to human participants, and collect judgements on the same 0-100 scale for all queries.
@@
Figure 5: Human game fun ratings correlate well with several distinct game features: entropy over game outcomes as predicted by participants themselves; outcome entropy predicted under our model; predicted advantage over a random agent given our model’s game play; expected game length under our model; and LLM-based estimates given the game specification. Far right plots show correlations in a combined regression model fit to all features predicted under our model (entropy, advantage, and length); and additionally incorporating the LLM-based estimates. Error bars depict standard error.
@@
People’s game outcome judgements are best predicted by partial, goal-directed game simulations. Subjects provided judgements estimating both the likelihood of a draw, and the likelihood that the first player would win if the game did not end in the draw, which collectively estimate the distribution and expected utility over all possible outcomes (draw, first player wins, second player wins) for any given game. Fig. 3 shows correlations between human estimations of game payoffs (expected utility, where $-1$ is a loss and $1$ is a win)—aggregated across all participants and game stimuli—and our model (Ours, blue) as well as the alternative baselines. As shown in Fig. 3, our model correlates well with human estimates ($R^{2}=0.852$, $p$ $<$ 5e-4).
@@
Notably, our model is a much better fit to human predictions than both the more naive models (such as the random partial simulation model, orange), and the more optimal models (both the depth-5 lookahead search, red, and MCS, pink). In particular, the game outcome predictions under the MCS baseline reveal that there is essentially no uncertainty about game outcomes under optimal play. That is, two true experts, or an agent estimating play from many more simulations, would expect to draw, or win based on their player ordering, every time. A qualitative analysis of human scratchpad data supports the conclusion that while most subjects spontaneously do seem to simulate games, they draw game estimates from very few games (almost no participants reset and then reused the scratchpad more than $5$ times), and imperfect game simulations (see examples in Fig. 2).
@@
Fig. 4 shows mean absolute error between the predicted payoff under models and human judgements on games, further split among the subcategories of our novel game variants. As seen in Fig. 4, our model (blue) is close (and usually closest) to human judgements within each of these subcategories. We also find that the LLM-based models (both the zero-shot and chain-of-thought baselines) are in fact relatively close to human judgements for the games that are the most obviously related to Tic-Tac-Toe (the $M$ in a row games on a square $N\times N$ board), but deviate much more from human judgements on seemingly minor variants—$M$ in a row games on rectangular boards or infinite boards—possibly because these games are relatively out-of-distribution compared to the more familiar variants found in background linguistic data.
@@
People’s game fun judgements can be predicted by their own and our model’s game outcome estimates. We evaluate several game-play-based features which we find each correlate with subject’s game fun estimates: 1) an estimate of whether games are fair and balanced based on the entropy over game outcomes—both predicted by humans themselves and predicted under our model; 2) an estimate of whether a game is reasonably challenging based on relative advantage between our model and a random agent (where expected utility is computed given random assignment of player order); and 3) expected game length. Interestingly, we also find the LLM-based CoT estimates of game fun to be well-correlated with human judgments. Fig. 5 summarizes these individual features, as well as fits from regression models fit to our model-derived features (and including the additional LLM-based prediction).

### v2

@@ Abstract
People can evaluate features of problems and their potential solutions well before we can effectively solve them. When considering a game we have never played, for instance, we might infer whether it is likely to be challenging, fair, or fun simply from hearing the game rules, prior to deciding whether to invest time in learning the game or trying to play it well. Many studies of game play have focused on optimality and expertise, characterizing how people and computational models play based on moderate to extensive search and after playing a game dozens (if not thousands or millions) of times. Here, we study how people reason about a range of simple but novel Connect-N style board games. We ask people to judge how fair and how fun the games are from very little experience: just thinking about the game for a minute or so, before they have ever actually played with anyone else, and we propose a resource-limited model that captures their judgments using only a small number of partial game simulations and almost no look-ahead search.
@@ Introduction
Figure 1: (A) Design of 121 grid games varying the game environment, dynamics, and win conditions. (B) Our intuitive game theory model simulates bounded game play under a fast but general agent model to draw probabilistic inferences about novel games.
The formal and empirical study of games often focuses on optimality and expertise. Game theory in mathematical economics, as studied by \@BBOPcite\@BAP\@BBNNash (1950)\@BBCP for example, largely seeks to characterize equilibrium states reached under rational play. Computational game playing models, from classic AI expert systems like Deep Blue to more recent neurally guided models like AlphaGo and AlphaZero \@BBOPcitep\@BAP\@BBN(Silver et al., 2016, 2017, 2018)\@BBCP, aim to match or exceed the abilities of human experts, generally by evaluating millions of simulated games. A rich body of related cognitive science research documents how human experts play after extensive experience on a given game \@BBOPcitep\@BAP\@BBN(Newell & Simon, 1972; Gobet et al., 2004)\@BBCP. Experts seem to learn highly specialized and game-specific representations, like memorized chess states and openings, that enable them to effectively search or approximate the results of searching much further ahead in a game than novice players \@BBOPcitep\@BAP\@BBN(Chase & Simon, 1973; van Opheusden et al., 2023)\@BBCP.
@@
The present study focuses on this higher-level capacity to flexibly evaluate novel inference and decision making problems, using a set of novel strategy games that vary the environment, dynamics, and win conditions of more familiar grid games as a test bed for general problem evaluation. We propose a computational model (Fig. 1b) that implements novice but goal-directed agents as fast but search-limited planners. Our model estimates the value of intermediate game state and move pairs based on simple but general concepts about game play, such as making progress towards a sparse overall goal (like making $M$ in a row) and preventing opponent progress when in a competitive setting. We then model how people evaluate games overall by nesting this agent model within a sample-based inference procedure that estimates game outcomes from a limited number of partial game simulations. While we focus here on games, the underlying building blocks of this “intuitive game theory” model are designed to capture more basic multi-agent planning and probabilistic inference components applicable to a much broader range of problems—where people are called on to make good guesses and good bets in novel situations, and to evaluate their prospects of success along with the emotional reward of engaging \@BBOPcitep\@BAP\@BBN(Allen et al., 2024)\@BBCP.
@@
We find that our intuitive game theory model, despite using only a single step of search in its agent model and a small number of partial game simulations to estimate game play, well predicts participants’ judgments of game outcomes on novel games. We compare our model against several alternative models, which either (i) consider even more naive game agent models, or conversely (ii) expend much more computational effort on search and game simulation to more optimally predict outcomes. We also compare against two language-model-based baselines with no explicit agent model that reason solely from background knowledge about other more familiar games.
Using our model, we also show that several game-related factors correlate with human fun ratings on these games: a fairness factor based on estimated entropy over game outcomes; a challengingness factor based on the estimated advantage our model has in playing against a random agent; and a fun factor based on language model predictions. Combined into a joint regression model, these factors together can predict most of the variance in human fun judgments.
Category Example(s) Count
@@
To choose any single action, an optimal agent model should estimate and then seek to maximize downstream utility with respect to all available actions from the current game state. Many general game play models estimate utility for novel games using expensive look-ahead search, ideally to terminal reward-generating game states \@BBOPcitep\@BAP\@BBN(Genesereth & Thielscher, 2014; Yannakakis & Togelius, 2018)\@BBCP. For instance, the class of games we consider here defines very sparse utility functions—winning or losing is only determined based on a final state in which a player or their opponent has placed $M$ pieces in a row. In cases where $M$ is large or players are playing on large boards (e.g., the goal is to make 8 in a row on a 20 by 20 board), this means searching over potentially long trajectories of game play to evaluate any single action.
To then draw probabilistic judgments about game outcomes (such as how likely a given non-deterministic player is to win on average), generic sample-based inference based on this agent model would require running multiple full game simulations, wherein each agent action would require downstream search to these terminal states.
In contrast, our intuitive game theory model aims to produce fast, probabilistic inferences over a broad range of novel games. This model simulates agent game play based on basic but general assumptions about goal-oriented players that can be estimated directly from intermediate states, then draws graded judgments based on partial game simulations over a range of game depths rather than always simulating to terminal states.
@@
The core of our model is a simulated agent that can play from game specifications. This agent model has a general search-limited way of assigning values over possible moves given current board states.
@@
Specifically, given an intermediate board state, our player model assigns an overall value score to each possible next move (each open position on the board) based on the following functions:
@@
    The maximum number of live connected (contiguous) pieces that the position entails on any of the allowed winning directions (horizontal, vertical, or diagonal in general, modulated with respect to any player-specific restrictions on win directions), $n_{1}\in\mathbb{Z}^{+}$. By live we mean connected pieces that have enough open positions on the corresponding direction such that they can be extended to form a winning $m$-in-a-row. For example, if $m=4$, then three connected pieces that are blocked by the opponent on both sides are not live. If $n_{1}$ equals the winning $m$, we add an additional 1 to $n_{1}$ to magnify the value of winning with respect to other states. This function captures the general intuition that players are goal-directed and try to make progress towards winning. That is, players derive utility from intermediate states towards the most rewarding terminal state. While there are many possible subgoals that a player could consider which might make partial progress towards terminal states (e.g., placing 3 in a row by placing two unconnected pieces and then placing one more piece in the center), we choose a simple, easy to calculate function that reflects a straightforward intuition (making $m$ in a row by previously building up to $m-1$ in a row, $m-2$ in a row, and so on via connected pieces). Variants of this function could be studied in future work, some of which could reflect more expert (rather than naive) play. For instance, we might reward positions based on the total number of live connected pieces that the position entails across all of the directions.
@@
    Relatedly, the maximum number of live connected contiguous pieces that the position blocks the opponent from having, $n_{2}\in\mathbb{Q}^{+}$, also considered with respect to the win-directions applicable to the opponent. This function captures the intuition that competitive players in antagonistic two-player games should also attempt to block opponent progress towards opponent goals. Note that this function is simple, symmetric to our definition of player progress (because players in our games have related and opposing goals). It also effectively evaluates just a single step of look-ahead search relative to a opponent who is also relatively naive (that is, an opponent who at the next step would simply attempt to make progress towards their nearest victory along a simple policy). We subtract $0.5$ from $n_{2}$ to reflect people’s tendency to weigh offense over defense \@BBOPcitep\@BAP\@BBN(Crowley & Siegler, 1993)\@BBCP, so blocking the opponent’s $\hat{n}$ in a row is not as good as making an $\hat{n}$ in a row for oneself (but is better than making an $\hat{n}-1$ in a row). However, if $n_{2}$ equals the winning $m$, we do not subtract $0.5$, similarly to upweight winning. One could consider other related or more complex formulations that also capture the notion of blocking opponent progress.
@@
Figure 3: Expected payoff across models against humans-predicted payoff. Each point represents the payoff for one of the $n=121$ game stimuli. Error bars in the y-axis are over individual human judgments per game and depict standard error. The $R^{2}$ values in the parentheses indicate 95% CI.
These components together reflect the intuitions that people often but not always place pieces around the center at the beginning, people naturally try to make progress towards winning, and people know they should secure winning moves and block the opponent from having those. The value of a position $p$ in board state $s$ is then given by a sum of exponential terms based on the components:
$\begin{split}V(p,s)=2^{(1-d)}+2^{n_{1}}+2^{n_{2}.}\end{split}$ (1)
We apply a softmax (temperature = $1$) to the set of values to obtain a probability distribution over possible moves in a state, and the agent samples a move from this distribution.
This agent model simulates game play—it captures intuitions about how a rational but relatively naive, resource-bounded player might choose an action from any given board state, on any of the games in our dataset. Given this model, how might someone draw graded inferences about the game itself, like about how fair the game is or how likely a player is to reach a particular outcome? We give a straightforward “game inference” model by nesting our agent model inside a sample-based probabilisitic inference model, which uses simulations of game play to estimate these kinds of features of a given game.
More specifically, for any given game, our model estimates inferences about any specific game query (e.g., what is the likelihood of any specific game outcome, win, loss, or draw; what is the likelihood that the first player will encounter any particular outcome?) by evaluating these queries with respect to $k$ sampled simulations of two players playing the game. This is a probabilistic estimator, as our agent model is stochastic. In keeping with our focus on fast, resource-limited reasoning about games, we specify two additional parameters that can bound the resources involved in inference. First, we estimate inference queries with a small number of game simulations (we use $k=20$ in our reported results), building upon the idea that people use a relatively small number of simulations to form beliefs and make decisions \@BBOPcitep\@BAP\@BBN(Vul et al., 2014; Icard, 2016)\@BBCP. Second, we allow our model to draw estimates using partial simulations: for each simulation, we randomly sample a bound on the total number of moves we will simulate (we use a uniform distribution over moves from $1$ to the size of the game board). We estimate the game outcome of any given simulation only based on the board state reached in that bounded set of moves (if a terminal win state hasn’t been reached, we simply score the game as a draw, though future work could use other estimators based on our heuristic board-state utility estimate). This bound reflects the notion that people might in fact draw estimates based on incomplete simulation of game play, a behavior qualitatively observed in the human scratchpad usage data.
This general inference algorithm, using a small number of partial, randomly bounded simulations, themselves simulated using a search-limited agent model, can be used to estimate many different game outcome queries. We focus on a standard one, expected payoff for the game (where we say winning incurs a reward of $+1$, losses incur $-1$, and draws are 00).
@@
Participants We recruit $484$ participants from Prolific \@BBOPcitep\@BAP\@BBN(Palan & Schitter, 2018)\@BBCP. Each participant was randomly presented with $10$ games sampled from our stimuli, as well as regular Tic-Tac-Toe (won by making 3 in a row on a $3\times 3$ board) to normalize game judgments. We collect approximately $20$ judgments per game stimuli for each game reasoning query. Participants were paid at a base rate of \$12.5/hr with an optional bonus up to \$15/hr; the full experiment approximately took $25$ minutes.
@@
Game outcome and game fun judgments Subjects were randomly assigned to one of two game reasoning conditions. In the game outcome evaluation condition, subjects produced judgments on a continuous 0-100 probability scale to predict the likelihood of a first player win (if the game does not end in a draw, assuming both players play reasonably, how likely is it that the first player is going to win (not draw)?) and a draw (assuming both players play reasonably, how likely is the game to end in a draw?). In the game fun condition, subjects instead assessed the likelihood that the game is fun (how fun is this game?) on a confidence scale spanning 0 (the least fun of this class of game) to 100 (the most fun of this class of game).
@@
Participants produced judgments about each game based on a linguistic game specification. We additionally provided participants with an interactive scratchpad board that they were told they could, but were not required, to use to inform their judgments; the scratchpad permitted automatically placing pieces of different colors (to simulate players) and could be cleared entirely to begin a new game. Participants were required to consider each game for at least $60$ seconds before producing game judgments; on average, participants took 87.2 ($\pm$ 2.5 STE) seconds per game in the game outcome condition and 83.4 ($\pm$ 2.41 STE) seconds per game in the fun rating condition.
@@
Figure 4: Absolute difference between human- and model-predicted payoff, broken down by game category. Error bars depict 95% CI over games within each game category (number of games in each game category is depicted in 1).
@@
Novel game creation After answering all game reasoning queries, we additionally asked participants to create a new grid-based game variant that they would find fun. Participants wrote a linguistic game specification, describing the board size and win conditions. As in the game judgment queries, participants were again provided an optional scratchpad and required to spend $60$ seconds before submitting a response. The scratchpad enabled participants to try out the game they intend to create. After specifying a game, participants were asked to answer the same game reasoning query (either game outcomes or game fun) about their own game.
Model game evaluations and alternative models
@@
We estimate simulated game evaluation judgments under our model. We also implement a range of computational models designed to span both weaker estimates of game play (using more naive agent models, or without any explicit agent models) and more optimal estimates of game play (using additional search and game simulations to terminal states).
Ours (Partial game simulations, 1-step heuristic over actions): Our model draws probabilistic game judgments via sample-based inference, where samples are partial game simulations rolled out under a subgoaling agent model that only estimates utilities over the immediate next actions according to equation 1. To match the sample size of the human experiment, we simulate $n=20$ judgments per game, where each judgment is estimated from $k=20$ samples.
@@
Partial game simulations, random actions: This baseline estimates outcomes using partial game simulations, but substitutes our sub-goaling agent model with a more naive (“random”) agent model that selects actions from a uniform distribution over valid moves. We again simulate $n=20$ judgments per game and $k=20$ samples per judgment.
@@
Full game simulations, random actions: This baselines infers outcomes from complete game simulations to terminal states, but using the random agent model. We again simulate $n=20$ judgments per game and $k=20$ samples per judgment.
[…]
@@
Figure 5: Human game fun ratings correlate well with several distinct game features: entropy over game outcomes as predicted by participants themselves; outcome entropy predicted under our model; predicted advantage over a random agent given our model’s game play; expected game length under our model; and LLM-based estimates given the game specification. Far right plots show correlations in a combined regression model fit to all features predicted under our model (entropy, advantage, and length); and additionally incorporating the LLM-based estimates. Error bars depict standard error. Numbers in parentheses indicate 95% CI.
Full game simulations with Monte Carlo Tree Search (MCTS) This baseline infers outcomes using full game simulations played using an MCTS-based agent model \@BBOPcitep\@BAP\@BBN(Genesereth & Thielscher, 2014)\@BBCP. We use a standard, upper-confidence-bound MCTS implementation; pseudocode can be found in the appendix (Algorithm 5). Unlike our main model or the approximate expert model, MCTS does not use game-specific heuristic features, and instead estimates intermediate utilities by expanding a search tree guided by repeated random rollouts to terminal states. MCTS algorithms are commonly used to approximate oracle or optimal gameplay in arbitrary games \@BBOPcitep\@BAP\@BBN(Genesereth & Thielscher, 2014)\@BBCP, as they are empirically efficient but accurate estimators of optimal play for many games (they effectively explore the search tree to recover optimal utility estimates, and converge to optimal estimators in the limit). We use MCTS as an approximate ‘oracle’ on our games and estimate actions using a large ($k=10,000$) number of tree expansions. As with the approximate expert model, due to computational costs, we estimate the expected payoffs using 50 simulations per game (2.5 simulations per subject) – but again, this model has less variance in outcome estimates per game simulation, as it tends to approximate more optimal play.
Large Language Model (LLM) game evaluations We also implement two LLM baselines to evaluate game reasoning with no explicit agent model, but rather, models which can draw on background knowledge about classic grid games and variants (including extensive information, for instance, about the optimal solutions to games like Tic-Tac-Toe and Connect-4). We use GPT-4 \@BBOPcitep\@BAP\@BBN(OpenAI, 2023)\@BBCP as our base LLM and implement two LLM-based game evaluation baselines: a zero-shot baseline (which is directly queried with the linguistic game specification and game reasoning query), and a zero-shot chain-of-thought (CoT) baseline \@BBOPcitep\@BAP\@BBN(Kojima et al., 2022; Wei et al., 2022)\@BBCP, which is first prompted to produce additional reasoning in text before answering a query (a standard method used to account for the computational limits in directly producing answers to questions given the LLM architecture). In both baselines, we prompt the model with a lightly-modified variant of the full experiment instructions shown to human participants, and collect judgments on the same 0-100 scale for all queries.
@@
People’s game outcome judgments are best predicted by partial, goal-directed game simulations. Subjects provided judgments estimating both the likelihood of a draw, and the likelihood that the first player would win if the game did not end in the draw, which collectively estimate the distribution and expected utility over all possible outcomes (draw, first player wins, second player wins) for any given game. Fig. 3 shows correlations between human estimations of game payoffs (expected utility, where -1 is a loss and 1 is a win)—aggregated across all participants and game stimuli—and our model (Ours, blue) as well as the alternative baselines. As shown in Fig. 3, our model correlates well with human estimates ($R^{2}=0.86$).
@@
Notably, our model is a much better fit to human predictions than both the more naive models (such as the random partial simulation model, orange), and the more optimal models (both the approximate expert, red, and MCTS, pink). In particular, the game outcome predictions under the approximate expert and the MCTS baselines reveal that under deeper search, there is much less uncertainty over game outcomes that does not match the gradedness and systematic deviations from optimality that humans demonstrate. A qualitative analysis of human scratchpad data supports the conclusion that while most subjects spontaneously do seem to simulate games, they draw game estimates from very few games (almost no participants reset and then reused the scratchpad more than $5$ times) and imperfect game simulations (see examples in Fig. 2).
@@
Fig. 4 shows mean absolute error between the predicted payoff under models and human judgments on games, further split among the subcategories of our novel game variants. As seen in Fig. 4, our model (blue) is close (and usually closest) to human judgments within arguably all but the last subcategories. We also find that the LLM-based models (both the zero-shot and chain-of-thought baselines) are in fact relatively close to human judgments for the games that are the most obviously related to Tic-Tac-Toe (the $M$ in a row games on a square $N\times N$ board), but deviate much more from human judgments on seemingly minor variants—$M$ in a row games on rectangular boards or infinite boards—possibly because these games are relatively out-of-distribution compared to the more familiar variants found in background linguistic data.
@@
People’s game fun judgments can be predicted by their own and our model’s game outcome estimates. We evaluate several game-play-based features which we find each correlate with subject’s game fun estimates: 1) an estimate of whether games are fair and balanced based on the entropy over game outcomes—both predicted by humans themselves and predicted under our model; 2) an estimate of whether a game is reasonably challenging based on relative advantage between our model and a random agent (where expected utility is computed given random assignment of player order) using full game simulations; and 3) expected game length under our model. Interestingly, we also find the LLM-based CoT estimates of game fun to be well-correlated with human judgments. Fig. 5 summarizes these individual features, as well as fits from regression models fit to our model-derived features (and including the additional LLM-based prediction). Better understanding what features of a game drive participants’ funness judgments are ripe grounds for future work.